OpenStudy (anonymous):

FAN MEDAL INSTAGRAM Which of the following options is an equivalent function to f(x) = 4(3)2x?

OpenStudy (anonymous):

F(x)=4*3* 2x f(x)=24x

OpenStudy (anonymous):

f(x) = 36x f(x) = 4(9)x f(x) = 144x f(x) = 4(6x)

OpenStudy (anonymous):

So D

OpenStudy (anonymous):

Yes it is D since 4*6x=24x

OpenStudy (anonymous):

Follow me at jherrera.29

OpenStudy (anonymous):

i agree with you